OpenStudy (anonymous):

What is the sum of the arithmetic sequence 3, 9, 15..., if there are 22 terms? 1,452 1,728 2.028 2,268

OpenStudy (anonymous):

\[ a_1=3\\ a_n= a_{n-1} +6 \]

OpenStudy (anonymous):

Here are the first 22 elements {3, 9, 15, 21, 27, 33, 39, 45, 51, 57, 63, 69, 75, 81, 87, 93, 99, \ 105, 111, 117, 123, 129} You can sum them or you can use the formula.

OpenStudy (anonymous):

Their sum is 1452 I will explain it to you very soon

OpenStudy (anonymous):

You can show that \[ a_n= 3 + 6(n-1),\, n =1, 2,\cdots 22 \]

OpenStudy (anonymous):

\[ \sum_{n=1}^{22} a_n=\sum_{n=1}^{22} (3 + 6(n-1)= 6 (22) + 6 \frac {(21)(22)}2=1452 \]